# Robust $\mathcal{H}_{\infty}$ control and worst-case search in constrained parametric space

### v4

@@ III-B Properties of the objective functions
In this section, we consider a generic transfer function written as the Linear Fractional Transformation (LFT) [54, Chapter 10] $T_{zw}(s,\delta)=\mathcal{F}_{u}(M(s),\Delta)$, as depicted in Figure 1, where $\Delta=\textrm{diag}(\delta_{i}I_{n_{i}})$ ($n_{i}$ is the number of repetitions of $\delta_{i}$). This representation will be used to characterize either the internal stability, or the $\mathcal{H}_{\infty}$ or $\mathcal{H}_{2}$ norm of the transfer from $w$ to $z$; practical problems e.g. (2) can also be addressed after an appropriate partitioning.
@@
to have no direct transfer from $\tilde{p}$ to $\tilde{q}$. This change of variable is the same as in [54, Section 12.3.4], with the uncertainty block instead of a feedback controller, and will be useful to study the $\mathcal{H}_{2}$ norm.
@@ III-B1 $\mathcal{H}_{2}$ norm (function $h_{2}$)
and its gradient is given in equation (18).
While the proof mainly follows previously established results cited thereafter, it is necessary to present the following equations to obtain the explicit gradient expression (18). The assumption $D_{22}=0$ is necessary to ensure finiteness of the $\mathcal{H}_{2}$ norm in the nominal case $\Delta=0$. The assumption $D_{12}=0$ or $D_{21}=0$ is done for the same reason for $\Delta\neq 0$. The differentiability of $h_{2}$ in $\mathcal{D}_{s}=\{\delta\in\mathcal{D},T_{zw}(s,\delta)\text{ is stable}\}$ was shown in [45, Section 3]; it was originally derived for a controller rather than an uncertainty block, but the validity of the result extends straightforwardly to the latter. Let us write the state-space matrices of the transfer from $w$ to $z$ after closing the loop between $p$ and $q$:
@@ III-B2 $\mathcal{H}_{\infty}$ norm (function $h_{\infty}$)
c) $h_{\infty}$ is differentiable in $\mathcal{D}_{s}$ when only one frequency is active and of multiplicity one; in this case, its gradient is given in equation (19).
@@
with $\star$ denoting the redheffer star product. The transfer functions $T_{zw}$, $T_{qw}$ and $T_{zp}$ correspond to the transfer between the inputs $p,w$ and outputs $q,z$ of the matrix $M(s)$ as in Figure 1, but after closing the loop with the block $\Delta$. If $\omega_{0}$ is the single active frequency, such that the multiplicity of $\bar{\sigma}(T_{zw}(j\omega_{0},\delta))$ is 1, $h_{\infty}$ is differentiable and, as shown in [3], its derivatives read:
@@
Where $p_{\omega_{0}}$ and $q_{\omega_{0}}$ are the normalized right and left singular vectors of $T_{zw}(j\omega_{0},\delta)$ associated with $\bar{\sigma}(T_{zw}(j\omega_{0},\delta))$ (which is equal to $h_{\infty}(\delta)$).
From the results of Section II and Proposition 5.b), we deduce that we can use smooth optimization, and in particular SQP (Algorithm 1), to solve Problem (13). In particular, let us define the direction $d_{\delta}$ such that its components are defined as in equation (19) (even when $h_{\infty}$ is not differentiable), i.e. the i-th component of $d_{\delta}$ reads:
@@
In contrast, minimizing the lower-$C^{1}$ function $h_{\infty}(\delta)$, i.e. minimizing the $\mathcal{H}_{\infty}$ norm, would be a more complex, genuinely nonsmooth problem. In particular, an unconstrained minimum may occur at a point where many subgradients are not zero (as a more simple example, consider $f:x\rightarrow|x|$, which is also lower-$C^{1}$). For this problem, nonsmooth techniques in the literature include multidirectionnal search [7], gradient sampling [18], or Clarke subdifferential-based [8, 15]. This is also pointed out in [3] as a fundamental difference between min-max and min-min (or max-max) programs, the former (e.g. minimizing the $\mathcal{H}_{\infty}$ norm) being genuinely non-smooth while the latter (e.g. maximizing the $\mathcal{H}_{\infty}$ norm or equivalently minimizing $h_{\infty}$) having the favorable properties demonstrated above.
@@ III-B3 Spectral abscissa (function $a$)
Proposition 6.a) was established in [3] following the earlier work of [14]. These references build on the study of [19] concerning the directional derivatives of the spectral abscissa. b) is also shown in [3], it simply follows from the fact that the spectral abscissa is the maximum of the real parts of the eigenvalues, which are differentiable when the eigenvalues are simple [41]. For c), the expression (20) is straightforward from the expression of the Clarke subdifferential given in [3].
@@
In this paper, as for $h_{\infty}$, we rely on the upper-$C^{1}$ property of $a$ of Proposition 6.b) and on the results of Section II, when active eigenvalues are simple. Our numerical experiments are satisfactory; however, the case of non-simple active eigenvalues is admittedly a limitation of using smooth optimization algorithms for the worst-case search in stability.
@@ III-C Global exploration, local exploitation
Problems (12), (13) and (14) are constrained optimization problems involving the nonsmooth functions $a$ and $h_{\infty}$, and the differentiable function $h_{2}$. As demonstrated in Section II, Sequential Quadratic Programming may be used, as it will converge to KKT points of the optimization problems under the upper-$C^{1}$ assumption. This constitutes the main theoretical contribution of the present paper. From a practical perspective, however, we additionally investigate the use of Particle Swarm Optimization or Monte-Carlo sampling to perform a global exploration of the parameter space, with SQP subsequently ensuring efficient local exploitation.
@@ III-C1 Sequential Quadratic Programming (SQP)
III-C2 Particle swarm optimization (PSO)
PSO is a heuristic global optimization method that handles nonsmooth functions. It only relies on the evaluation of the function and does not depend on (sub)gradient information or on the regularity properties discussed in Section III-B; thus, it may help circumvent limitations encountered when the spectral abscissa fails to be upper-$C^{1}$, or even Clarke subdifferentiable. The nonlinear constraint is treated with a barrier method to penalize non admissible solutions: we reformulate problem (13) as:
$\begin{array}[]{lll}&\underset{\delta\in\mathbb{R}^{k}}{\text{minimize}}&h_{\infty}(\delta)+\tau\sum_{i}\max(0,c_{i}(\delta))\\ &\text{subject to }&-1\leq\delta_{i}\leq 1\end{array}$
and problems (12) and (14) can be treated similarly. The parameter $\tau$ can be set to a high value, and iteratively increased if necessary until a solution $\delta^{*}$ verifying all constraints $c_{i}(\delta^{*})\leq 0$ is found. The bounds on the $\delta_{i}$ are enforced by projecting the particles positions into the admissible space at each iteration. In our numerical testings, we used MATLAB routine pso described in [42]; PSO was used as a method to explore the parametric space (and the choice of parameters was done accordingly, as discussed in Section V-C), with local exploitation being performed by a subsequent SQP.
III-C3 Monte-Carlo (MC) sampling
Monte Carlo sampling is applied as follows: randomly pick a sample $\delta$ in the unit hypercube (considering for example uniform distributions for all parameters); evaluate $c(\delta)$: if positive, discard $\delta$, otherwise, evaluate the function $a(\delta)$ (respectively $h_{\infty}(\delta)$ or $h_{2}(\delta)$); repeat until having $N$ function evaluations. The worst-case $\delta^{*}$ corresponds to the lowest value of $a$ (respectively $h_{\infty}$ or $h_{2}(\delta)$). In our testings, Monte Carlo will also be followed by SQP for local exploitation.
In such a worst-case search problem, the number $N$ of function evaluations can be chosen to verify:
@@ III-C3 Monte-Carlo (MC) sampling
for an objective function $f$ (in our case: $a(\delta)$, $h_{\infty}(\delta)$ or $h_{2}(\delta)$). Note that this bound is independent from the number of uncertain parameters, from their variation ranges or distributions. Since this paper addresses deterministic worst-case search, it is natural to draw parameters from uniform distributions; note, however, that removing some inadmissible configurations implicitly modifies these distributions.
@@ IV-A Robust controller optimization algorithm
The Algorithm 2, described in a form that solves the standard $\mathcal{H}_{\infty}$ control problem (1) – the multi-objectives formulation is tackled in Section IV-B – follows the approach proposed in [3] that relies on iteratively constructing a subset $\mathcal{D}_{a}\subset\mathcal{D}$ of so-called active configurations as follows.
Step 1 optimizes the controller for all current active configurations. It uses the nonsmooth optimization algorithm of [8] implemented in the routine systune. Step 2 addresses the worst-case search in stability as described in Section III; if a worst-case configuration, i.e. a Karush-Kuhn-Tucker (KKT) point of the optimization problem, is found that leads to a spectral abscissa greater than a tolerance $\alpha_{max}$, it is added to $\mathcal{D}_{a}$. It is also possible to identify multiple KKT points during Step 2. Similarly, Step 3 tackles the worst-case search in performance and adds to $\mathcal{D}_{a}$ the worst-case configuration(s) that degrade the performance by more than a tolerance $\epsilon$. Finally, Step 4 performs a more exhaustive worst-case search, meaning that the optimization parameters are tuned to increase coverage of the parameter space and ensure a more thorough exploration of potential worst-case configurations, while in Steps 2 and 3 they are chosen to prioritize fast convergence to KKT points; additional quantitative details are provided in the application of Section V-D. For Steps 2, 3 and 4, the global exploration may be performed either with Particle Swarm Optimization (PSO), or Monte-Carlo (MC) if stochastic guarantees are sought; in both cases, a subsequent Sequential Quadratic Programming (SQP) refines the search until finding a KKT point.
Algorithm 2 Robust control optimization with constrained uncertainties
@@
Identify one or multiple KKT point(s) for the problem: $\delta^{*}=\arg\;\max_{\delta\in\mathcal{D}}\;\alpha(A(K^{*},\delta))$ using PSO/MC+SQP.
@@
Perform exhaustive worst-case search in stability and performance.
@@ IV-B Multi-objective formulation
In our numerical implementation, we used the expression on the left side of equation (22); indeed, it was faster to compute the norm of each requirement, than the norm of the block diagonal system.
More generally, we consider the multi-objectives problem (2). As explained in [4], the controller optimization of Step 1 treats it as :
@@
i.e. the algorithm has converged when the worst-case search degrades the soft requirement by less than $1+\epsilon_{1}$, and the hard requirements are less than $1+\epsilon_{2}$. Finally, we point out that, in our implementation, the PSO computes all norms involved in equation (23) and determines their maximum; whereas the subsequent SQP only evaluates the transfer function that achieved this maximum.
V Application: a large flexible spacecraft
@@ V-A Plant model and uncertain parameters
The benchmark used in this study is designed to be challenging regarding robustness. We consider a spacecraft composed of one rigid body connected to two large flexible solar panels. The clamped-free natural frequencies are between 0.20 and 0.85 rad/s for each panel. The resulting 8 nominal free-free frequencies are between 0.20 and 4.2 rad/s for the whole spacecraft. Uncertainties are considered on various mechanical properties:
  • Each clamped-free frequency exhibits a common 10% across both panels, along with an additional 5% asymmetric uncertainty. Modal damping is 0.001 for each clamped-free mode.
  • The mass, center of mass, inertia (MCI) properties of the solar panel are uncertain, as well as the modal participation factors of the flexible modes.
  • The angle between the central body and the solar arrays may have any value between $\pi$ and $\pi$.
  • The central body also has uncertain MCI properties.
  • We also consider an uncertain stiffness and damping in the rotation of the solar array, as a simple solar array driving mechanism (SADM) model.
In total, there are 43 uncertain parameters $\delta_{i}$ (normalized such that $1\leq\delta_{i}\leq 1$, without loss of generality) regrouped in a vector $\delta$. Noting $M_{P}(\delta)$ the mass matrix at attachment point $P$ of a solar panel, and $L_{P}(\delta)$ the matrix of modal participation factors, the robustness analysis should exclude any configuration $\delta$ such that the residual mass matrix defined in equation (24) is not positive definite (indeed, such configurations do not represent a physical system). Hence the nonlinear constraint:
@@
The spacecraft model $G(s,\delta)$ is built with the multibody approach detailed in [1] and implemented in the Satellite Dynamics Toolbox Library (SDTlib) [2, 48]. The block $\Delta_{G}$ of the Linear Fractional Representation (LFR) model, such that $G(s,\delta)=\mathcal{F}_{u}(M_{G}(s),\Delta_{G})$ (upper LFT), has size 204. We emphasize that this model contains all parametric configurations in a continuous manner, defined at each subsystem level based on the physics equations without any interpolation or other approximation (besides the linearity assumption). The LFR model of the spacecraft includes all configurations for $1\leq\delta_{i}\leq 1$ even though only $\sim 22\%$ of them respect the nonlinear constraint (25).
@@
The constraint (25) is not differentiable everywhere, which contradicts the assumptions of Section 2. However, since this is the formulation that naturally arises in our application, we still test it numerically.
Remark 9.
In addition to providing the analytical expressions of the gradients given in Section III, the LFR representation also enables fast model evaluation at given parametric configuration by simply substituting the $\Delta_{G}$ block.

### v5

@@ III-B Properties of the objective functions
In this section, we consider a generic transfer function written as the Linear Fractional Transformation (LFT) [41, Chapter 10] $T_{zw}(s,\delta)=\mathcal{F}_{u}(M(s),\Delta)$, as depicted in Fig. 1, where $\Delta=\textrm{diag}(\delta_{i}I_{n_{i}})$ ($n_{i}$ is the number of repetitions of $\delta_{i}$). This representation will be used to characterize either the internal stability, or the $\mathcal{H}_{\infty}$ or $\mathcal{H}_{2}$ norm of the transfer from $w$ to $z$; practical problems e.g. (2) can also be addressed after an appropriate partitioning.
@@
to have no direct transfer from $\tilde{p}$ to $\tilde{q}$. This change of variable is the same as in [41, Section 12.3.4], with the uncertainty block instead of a feedback controller, and will be useful to study the $\mathcal{H}_{2}$ norm.
@@ III-B1 $\mathcal{H}_{2}$ norm (function $h_{2}$)
and its gradient is given in Eq. (18).
While the proof mainly follows previously established results cited thereafter, it is necessary to present the following equations to obtain the explicit gradient expression (18). The assumption $D_{22}=0$ is necessary to ensure finiteness of the $\mathcal{H}_{2}$ norm in the nominal case $\Delta=0$. The assumption $D_{12}=0$ or $D_{21}=0$ is done for the same reason for $\Delta\neq 0$. The differentiability of $h_{2}$ in $\mathcal{D}_{s}=\{\delta\in\mathcal{D},T_{zw}(s,\delta)\text{ is stable}\}$ was shown in [34, Section 3]; it was originally derived for a controller rather than an uncertainty block, but the validity of the result extends straightforwardly to the latter. Let us write the state-space matrices of the transfer from $w$ to $z$ after closing the loop between $p$ and $q$:
@@ III-B2 $\mathcal{H}_{\infty}$ norm (function $h_{\infty}$)
c) $h_{\infty}$ is differentiable in $\mathcal{D}_{s}$ when only one frequency is active and of multiplicity one; in this case, its gradient is given in Eq. (19).
@@
with $\star$ denoting the redheffer star product. The transfer functions $T_{zw}$, $T_{qw}$ and $T_{zp}$ correspond to the transfer between the inputs $p,w$ and outputs $q,z$ of the matrix $M(s)$ as in Fig. 1, but after closing the loop with the block $\Delta$. If $\omega_{0}$ is the single active frequency, such that the multiplicity of $\bar{\sigma}(T_{zw}(j\omega_{0},\delta))$ is 1, $h_{\infty}$ is differentiable and its derivatives read:
@@
where $p_{\omega_{0}}$ and $q_{\omega_{0}}$ are the normalized right and left singular vectors of $T_{zw}(j\omega_{0},\delta)$ associated with $\bar{\sigma}(T_{zw}(j\omega_{0},\delta))$ [3].
From the results of Section II and Proposition 5.b), we deduce that we can use SQP (Algorithm 1) to solve Problem (13). In particular, let us define the direction $d_{\delta}$ such that its components are defined as in Eq. (19) (even when $h_{\infty}$ is not differentiable), i.e. the i-th component of $d_{\delta}$ reads:
@@
In contrast, minimizing the lower-$C^{1}$ function $h_{\infty}(\delta)$, i.e. minimizing the $\mathcal{H}_{\infty}$ norm, would be a more complex, genuinely nonsmooth problem. In particular, an unconstrained minimum may occur at a point where many subgradients are not zero (as a more simple example, consider $f:x\rightarrow|x|$, which is also lower-$C^{1}$). Moreover, descent directions are not provided by any subgradient. For this problem, nonsmooth techniques in the literature include multidirectionnal search [7], gradient sampling [18], or Clarke subdifferential-based [8, 15]. This is also pointed out in [3] as a fundamental difference between min-max and min-min (or max-max) programs, the former (e.g. minimizing the $\mathcal{H}_{\infty}$ norm) being genuinely non-smooth while the latter (e.g. maximizing the $\mathcal{H}_{\infty}$ norm or equivalently minimizing $h_{\infty}$) having the favorable properties demonstrated above.
@@ III-B3 Spectral abscissa (function $a$)
Proposition 6.a) was established in [3] following the earlier work of [14]. These references build on the study of [19] concerning the directional derivatives of the spectral abscissa. b) is also shown in [3], and simply follows from the fact that the spectral abscissa is the maximum of the real parts of the eigenvalues, which are differentiable when the eigenvalues are simple [32]. For c), the expression (20) is straightforward from the expression of the Clarke subdifferential given in [3].
@@
In this paper, as for $h_{\infty}$, we rely on the upper-$C^{1}$ property of $a$ of Proposition 6.b) and on the results of Section II, when active eigenvalues are simple. However, the case of non-simple active eigenvalues is admittedly a limitation of using smooth optimization algorithms for the worst-case search in stability.
@@ III-C Global exploration, local exploitation
Problems (12), (13) and (14) are constrained optimization problems involving the nonsmooth functions $a$ and $h_{\infty}$, and the differentiable function $h_{2}$. As demonstrated in Section II, Sequential Quadratic Programming may be used, as it will converge to KKT points of the optimization problems under the upper-$C^{1}$ assumption. This constitutes the main theoretical contribution of the present paper. From a practical perspective, however, we additionally investigate the use of Monte-Carlo sampling to perform a global exploration of the parameter space, with SQP subsequently ensuring efficient local exploitation.
@@ III-C1 Sequential Quadratic Programming (SQP)
III-C2 Monte-Carlo (MC) sampling
Monte Carlo sampling is applied as follows: randomly pick a sample $\delta$ in the unit hypercube; evaluate $c(\delta)$: if positive, discard $\delta$, otherwise, evaluate the function $a(\delta)$ (respectively $h_{\infty}(\delta)$ or $h_{2}(\delta)$); repeat until having $N$ function evaluations. Since this paper addresses deterministic worst-case search, it is natural to draw parameters from uniform distributions; note, however, that removing some inadmissible configurations implicitly modifies these distributions. To leverage the local search enabled by SQP, we envision three ways to perform the Monte-Carlo sampling.
First, assume a probability $p$ that a random sample lies in the basin of attraction of the global optimum, i.e. that a gradient-based local optimization starting from this sample converges to the global optimum. If $N$ samples are drawn and all serve as initial point for an SQP run, the probability of finding the global optimum is $1-(1-p)^{N}$. Thus, given a target probability $p$ (reflecting the size of the basin of attraction of the global optimum that we wish to be able to detect), and a confidence level $P$ such that $1-(1-p)^{N}\geq P$, the required number $N$ of samples is
$N\geq\frac{\mathrm{ln}(1-P)}{\mathrm{ln}(1-p)}\;.$
A second possibility, in worst-case search problems, is to pick the number $N$ of function evaluations to verify:
@@ III-C2 Monte-Carlo (MC) sampling
where $\delta^{*}$ corresponds to the lowest value of the objective function $f$ (in our case: $a(\delta)$, $h_{\infty}(\delta)$ or $h_{2}(\delta)$) among the $N$ samples. Note that this bound is independent from the number of uncertain parameters, from their variation ranges or distributions. Then, a local search with SQP can be performed starting from this worst case.
A third option combines the two previous approaches. It consists of drawing $n$ groups of $N$ samples, and for each group, selecting the worst-case sample, denoted $\delta^{*}_{1},\delta^{*}_{2},...\delta^{*}_{n}$, to initialize a SQP search. Each of the $\delta^{*}_{i}$ verifies the probability of Eq. (21). Additionally, the minimum among all $\delta^{*}_{i}$ also verifies a condition of the same kind, but among $nN$ samples. This strategy enables the local search to be initialized from multiple points, thereby increasing the coverage of the parametric space compared to the second approach, while requiring fewer SQP runs (more costly than evaluating random points) than the first approach. This approach is applied in our numerical application in Section V-C.
@@ IV-A Robust controller optimization algorithm
The Algorithm 2, described in a form that solves the standard $\mathcal{H}_{\infty}$ control problem (1), follows the approach proposed in [3] that relies on iteratively constructing a subset $\mathcal{D}_{a}\subset\mathcal{D}$ of so-called active configurations as follows. The multi-objectives formulation is tackled in Section IV-B.
The set of active configurations is initialized with the nominal configuration. Step 1 optimizes the controller for all current active configurations, based on the nonsmooth optimization algorithm of [8] implemented in the routine systune. Step 2 addresses the worst-case search in stability as described in Section III. If a Karush-Kuhn-Tucker (KKT) point of the optimization problem is found that leads to a spectral abscissa greater than a tolerance $\alpha_{max}$, it is added to $\mathcal{D}_{a}$. It is also possible to identify multiple KKT points during Step 2. Similarly, Step 3 tackles the worst-case search in performance and adds to $\mathcal{D}_{a}$ the worst-case configuration(s) that degrade the performance by more than a tolerance $\epsilon$. Finally, Step 4 seeks to validate the design with a more exhaustive worst-case search, meaning that the optimization parameters are tuned to ensure a more thorough exploration of potential worst-case configurations. In contrast, in Steps 2 and 3, the parameters are chosen to prioritize fast convergence to KKT points; additional quantitative details are provided in the application of Section V-D. For Steps 2, 3 and 4, a global exploration is performed with Monte-Carlo (MC), and a subsequent Sequential Quadratic Programming (SQP) refines the search until finding a KKT point.
Algorithm 2 Robust control optimization with constrained uncertainties (largely inspired from [3])
@@
Identify one or multiple KKT point(s) for the problem: $\delta^{*}=\arg\;\max_{\delta\in\mathcal{D}}\;\alpha(A(K^{*},\delta))$ using MC+SQP.
@@
Perform exhaustive worst-case search in stability and performance using MC+SQP.
@@ IV-B Multi-objective formulation
In our numerical implementation, in this case, we used the expression on the left side of Eq. (22); indeed, it was faster to compute the norm of each requirement, than the norm of the block diagonal system.
More generally, we consider the multi-objectives problem (2). As explained in [4], the controller optimization [8] of Step 1 treats it as :
@@
i.e. the algorithm has converged when the worst-case search degrades the soft requirement by less than $1+\epsilon_{1}$, and the hard requirements are less than $1+\epsilon_{2}$. Finally, we point out that, in our implementation, the Monte-Carlo sampling computes all norms involved in Eq. (23) and determines their maximum; whereas the subsequent SQP only evaluates the transfer function that achieved this maximum.
V Illustration: a large flexible spacecraft
@@ V-A Plant model and uncertain parameters
The benchmark used in this study is designed to be challenging regarding robustness. We consider a spacecraft composed of one rigid body connected to two large flexible solar panels of clamped-free natural frequencies between 0.20 and 0.85 rad/s. Uncertainties are considered on various mechanical properties; in total, there are 43 uncertain parameters $\delta_{i}$ (normalized such that $1\leq\delta_{i}\leq 1$, without loss of generality) regrouped in a vector $\delta$. Noting $M_{P}(\delta)$ the mass matrix at attachment point $P$ of a solar panel, and $L_{P}(\delta)$ the matrix of modal participation factors, the robustness analysis should exclude any configuration $\delta$ such that the residual mass matrix defined in Eq. (24) is not positive definite (indeed, such configurations do not represent a physical system). Hence the nonlinear constraint:
@@
The spacecraft model $G(s,\delta)$ is generated with the multibody approach detailed in [1] and implemented in the Satellite Dynamics Toolbox Library [2, 37]. The block $\Delta_{G}$ of the Linear Fractional Representation (LFR) model, such that $G(s,\delta)=\mathcal{F}_{u}(M_{G}(s),\Delta_{G})$ (upper LFT), has size 204. The LFR model of the spacecraft includes all configurations for $1\leq\delta_{i}\leq 1$ even though only $\sim 22\%$ of them respect the nonlinear constraint (25).
@@
The constraint (25) is not guaranteed to be differentiable everywhere, which contradicts the assumptions of Section 2. However, since this is the formulation that naturally arises in our application, we still test it numerically.